A 142 g ball is tied to a string. It is pulled to an angle of 4.0 ∘ and released to swing as a pendulum. A student with a stopwatch finds that 12 oscillations take 13 s .

How long is the string?
Answer #1

time for 1 oscillation,
T = 13/12 s

use:
T = 2*pi*sqrt(L/g)
13/12 = 2*pi*sqrt (L/9.8)
L / 9.8 = 0.02973
L = 0.29 m
   = 29 cm
Answer: 29 cm